The horn of a stationary automobile is sounding a note of 580Hz. What 

frequency will you hear if you are driving towards this automobile at 

80km/hr? What change of frequency will you notice after you have passed?( 

speed of sound in air is 340m/s)


1
Expert's answer
2021-03-23T11:11:21-0400

"80\ km/hr=22.22\ m/s"


"(1)\ \ \nu=\nu_0(1+v/v_0)=580\cdot(1+22.22/340)=617.9\ (Hz)"


after you have passed


"(2)\ \ \nu=\nu_0(1-v/v_0)=580\cdot(1-22.22/340)=542\ (Hz)"


"\Delta\nu=580-542=38\ (Hz)" in relation to the frequency of the sound source.